# A spring–block theory of feature learning in deep neural networks

### v2

@@ Abstract
Deep learning theories are primarily built bottom-up, by analyzing the collective behavior of the myriad artificial neurons interacting in a data-driven energy from first principles. But fields like physics, biology, neuroscience, and economics, which have long dealt with comparable complexity (or, sometimes, worse) have long benefitted from both bottom-up and top-down, phenomenological approaches. We think that deep learning can similarly benefit from both paradigms. Our study attempts to introduce a simple phenomenological model to help understand feature learning—an essential aspect of modern real-world DNNs which requires us to simultaneously consider depth, nonlinearity, and noise, something that at the moment seems to be out of reach of the bottom-up paradigm. It is a priori not obvious that such a model exists: it requires identifying a suitable low-dimensional description of the system and demonstrating its approximately autonomous behavior. Inspired by a series of works on neural collapse and data separation, our theory elucidates the role of nonlinearity and randomness and suggests exciting connections with generalization and conjectures that are worth pursuing rigorously. Analogies such as spring-block models have historically played important roles across science; a prime example is the Burridge–Knopoff model in seismology [65] and related ideas in neuroscience [66]. We find it intriguing that a simple phenomenological model holds significant explanatory power for realistic DNNs, serving as a valuable tool for applying intuitive physical insights to complex, abstract objects. In closing, we mention that we initially considered various cascade structures from daily life as possible analogies to DNNs. For amusing videos involving a folding ruler, see the SM [57].
We are grateful to Hangfeng He, Weijie Su and Qing Qu for inspiring discussions about the law of data separation and training dynamics. Cheng Shi and Ivan Dokmanić were supported by the European Research Council Starting Grant 852821—SWING. Liming Pan acknowledges support from National Natural Science Foundation of China (NSFC) under Grant No. 62006122 and 42230406
@@
I S1. A folding ruler experiment
We first describe an experiment where a different cascaded mechanical system, a folding ruler, exhibits a phenomenology which is in some ways reminiscent of feature learning dynamics in DNNs. We emphasize that this reminiscence is anecdotal but nonetheless mention it since it motivated our work. The goal of the experiment is to show that noise can renegotiate the imbalance caused by friction. As shown in Fig. S1, we pin the left end of the folding ruler and pull the right end by hand. Due to friction, if we pull it very slowly and steadily, the outer layer extends far while the inner layers are close to stuck. This reminds us of lazy training where the outer layers take the largest proportion of the load. Conversely, shaking while pullings helps “activate” the inner layers and redistribute the force, and ultimately results in a uniformly distributed extension of each layer. Videos can be found at https://github.com/DaDaCheng/DNN_Spring
@@
II S2. Additional experimental results on various networks and datasets
@@
Similarly to MLPs, deep convolution neural networks (CNNs) also learn features and progressively collapse data. Here we explore how the observation that a uniform load distribution correlates with better generalization may help guide CNN training. In Fig. S2, we illustrate an experiment where we first train a CNN using Adam [67] at a learning rate of $10^{-4}$ on MNIST. The training nearly converges after $20\times 200$ epochs, resulting in a concave load curve (purple). Next, we introduce a $5\%$ dropout which causes the training to resume and the load curve to become linear (blue). Importantly, it also improves accuracy. Stronger noise (a $30\%$ dropout at $40\times 200$ epochs) at once results in a convex load curve and worse generalization.
Figure S2: The dynamics of load curves for a deep CNN. (a) test accuracy versus training loss. (b) the corresponding load curves during training. In the experiments, we introduce $5\%$ at epoch $20\times 200$ and $30\%$ dropout at epoch $40\times 200$.
In Fig. S3, we illustrate the in-class means of the $\ell$-th mid-feature $(\bar{\bm{{x}}}_{\ell}^{k})$ at epochs 4000, 8000, and 12000. The corresponding load curves at these times are concave, linear, and convex, respectively, as shown in Fig. S2.
Figure S3: In-class means of the mid layer features in deep CNN whose training dynamics is showed in Fig. S2. We plot first channel at each layer for digits 0, 4, and 9 at epochs 4000 (concave load curve), 8000 (linear load curve), and 12000 (convex load curve).
@@
This work primarily investigates the concavity and convexity of the training load curve. Notably, the test load curve demonstrates behavior similar to that of the training load curve. In Fig. S4 we show the load curves for both the training samples (solid line) and test samples (dashed line) on the MNIST dataset. We can see that with a small learning rate both both curves are concave. Increasing the learning rate makes them both linear; increasing it further makes them convex. The most linear curve corresponds to the highest test accuracy.
@@
The phenomenon that the highest accuracy occurs when the load curve is linear is universally observed across a wide range of parameters and datasets. It occurs, for example, for different depths in Fig. S5 and different widths in Fig. S6.
@@
It also occurs with other datasets and optimizers. We experiment with the FashionMNIST dataset in Fig. S7 with the same default setting in (a) and a different optimizer in (b); further experiments on CIFAR10 are shown in Fig. S8. The original color images are first resampled to a size of $3\times 10\times 10$ for computational reasons and then vectorized as input to an MLP. We note that $D_{0}$ (the data separation metric in the input layer) slightly deviates from linearity, likely due to the additional complexity of CIFAR10, but the subsequent layers closely follow the regular phenomenology discussed in the main text (Fig. S8(b)).
@@
The noise vs. nonlinearity phase diagram remains valid even when the data becomes more complex, as does the correlation between generalization and linearity of separation across layers when noise is introduced via dropout. A small deviation from this occurs when dropout is replaced by learning rate on CIFAR10: in this case the best generalization corresponds to slightly concave load curves (Fig. S9). Since learning rate is qualitatively very different from standard “noise” we expect that at some point it exhibits a refined phenomenology; this particular example is an intriguing avenue for further exploration.
@@
III S3. Details of numerical experiments and reproducibility
@@
In Figs. 1, 4, S4 S5, S6, S7, S8, S9, S10 and S12, the networks are $8$-layer fully connected MLPs (7 hidden layers), with layer width equal to $100$ unless stated otherwise. We use ReLU activations, BatchNorm in each layer, and no dropout as the default setting. All parameters are initialized using the default settings in PyTorch. Unless otherwise specified, the networks are trained using the ADAM optimizer on 2560 training samples from the MNIST dataset, with cross-entropy loss for classification tasks. The learning rate is set to 0.001 and the batch size is 2560 unless otherwise stated. The test accuracy is measured on the entire 10,000-sample test dataset after training 100 epochs. The results in all phase diagrams are averaged over 10 independent runs. Notably, when we replace BatchNorm with the scaling constant $c_{\ell}$ mentioned in the main text, we still observe similar convex–concave patterns in the phase diagram corresponding to noise and nonlinearity. However, without the adaptive BatchNorm, the load curves exhibit more fluctuations and are less smooth, and the variations between independent runs become more pronounced.
@@
In the left column of Fig. 4, we train the described DNN with learning rate of 0.0001 and batch size of 200 in panel (b); learning rate of 0.001, dropout of 0.1 and batch size of 100 in panel (c); learning rate of 0.003, and batch size of 50, and dropout of 0.2 in panel (d). In the right column we set $k=1,y=1$ for all three cases, and $\mu_{\rightarrow}=0.11,\mu_{\leftarrow}=0.03$ and $\varepsilon=0.006$ for (b); $\mu_{\rightarrow}=\mu_{\leftarrow}=0.04$ and $\varepsilon=0.01$ for (c); $\mu_{\rightarrow}=0.04,\mu_{\leftarrow}=0.12$ and $\varepsilon=0.011$ for (d). These parameters are chosen to produce qualitatively similar curves with the previous three characteristic training dynamics. For the spring experiments in Fig. 4, we use the same noise for all blocks $\xi_{\ell}(t)=\xi_{\ell^{\prime}}(t)$, to mimic the training of DNNs in which the randomness (e.g., data noise, learning rate, and batch size) in each layer are not independent. This synchronous noise results in more similar dynamics over epochs (in particular the fluctuations) but ultimately leads to similar load curves as independent noise as shown in Fig. 2.
@@
In Fig. 5(a), we adopt the setting from [41]. We use SGD (instead of ADAM) with learning rate of 0.001 to train a linear DNN. All weight matrices are initialized as random orthogonal matrices; the data is also a random orthogonal matrix with random binary labels. There is no batch normalization. We set the learning rate to $0.01$ to generate the “large stepsize” result (blue curve) and apply $10\%$ dropout to generate the “dropout” results (green curve).
@@
In Fig. S2, we consider a CNN with 16 channels and 7 convolutional layers on the same MNIST dataset. We use ReLU activations and BatchNorm between the convolutional layers. Pooling and upsampling are applied after each activation in such a way that each intermediate layer has the same shape, and a linear layer is applied after the final convolutional layer.
@@
IV S4. Some other metrics for data separation of intermediate features
@@
$\hat{D}_{\ell}:=\log\left(\mathrm{Tr}\left(\Sigma_{\ell}^{\mathrm{w}}{\Sigma_{\ell}^{\mathrm{b}}}^{+}\right)\right),$ (9)
@@
Figure S10: Phase diagram of the DNN load curve on the MNIST for nonlinearity vs. dropout with the data separation as $\hat{D}_{\ell}$ in Eq. (9).
The behavior of the two metrics notably diverges for linear networks where $\hat{D}_{\ell}$ cannot capture the differences in features across different layers (Fig. S11 (c) and (d)).
Figure S11: Comparison of data separation metrics $D_{\ell}$ (Eq. (3)) in (a, c) and $\hat{D}_{\ell}$ (Eq. (9)) in (b, d). The top two plots (a, b) show results for the ReLU DNN, while the bottom two plots (c, d) show results for the linear DNN. The curves in all four plots represent the same DNNs trained on MNIST with different dropout (dp) rates, as indicated in the legend in (c).
V S5. Regression
@@
$D_{\ell}:=\sqrt{\min_{\bm{{w}},\bm{{b}}}\frac{1}{2N}\sum_{(\bm{{x}}_{\ell},\bm{{y}})}\|\bm{{y}}-\bm{{w}}^{\top}\bm{{x}}_{\ell}+\bm{{b}}\|^{2}_{2}},$
where $N$ is the number of data pairs and the summation is taken over all data. As shown in Fig. S12, noisier training results in a convex load curve, whereas lazy training results in a concave curve.
Figure S12: Regression load curve on Diabetes dataset [68] for a 7 hidden-layer ReLU network. (i) learning rate of 0.001 without dropout. (ii) learning rate of 0.01 without dropout (iii) learning rate of 0.02 and with 30% dropout.
VI S6. Additional details about the spring block systems
VI.1 Friction in the second order system
@@
VI.2 Noisy equilibiria and separation of time scales
@@
VI.3 Second order Langevin equation
We can obtain a more standard Langevin dynamics formulation of Eq. (10) by adding noise to the velocity independently of the friction, i.e., computing the friction before adding noise. We note that this is less realistic from a physical point of view. The position of the $\ell$th block $x_{\ell}=\sum_{i=1}^{\ell}d_{i}$ then obeys the equation of motion $\mathrm{d}x=v\mathrm{d}t$ with
$\mathrm{d}v_{\ell}=\left(k(\bm{L}x)_{\ell}-f(v_{\ell})-\gamma v_{\ell}\right)\mathrm{d}t+\varepsilon(t)\mathrm{d}W_{\ell}(t),$ (12)

### v4

@@ Abstract
Deep learning theories are mostly built bottom-up, but fields like physics, biology, neuroscience and economics benefit from both bottom-up and top-down, phenomenological approaches. Mechanical analogies such as spring-block models play an important role across science; a prime example is the Burridge–Knopoff model in seismology [64] and related ideas in neuroscience [65]. We think that deep learning can similarly benefit from both paradigms, especially with complex phenomena like feature learning which require to simultaneously consider depth, nonlinearity, noise and data.
Our phenomenological model elucidates the role of nonlinearity and randomness and suggests exciting connections with generalization that may inspire new theory, but also, importantly, new approaches to hyperparameter tuning and model selection. The current theory applies to DNNs where inner layers are alike; extensions to heterogeneous DNNs will require refined definitions of data separation and new mechanical models.
We finally mention that we considered various cascade structures as possible analogies to DNNs; the SM links to real experiments with a folding ruler [53].
Figure 6: The dynamics of load curves for a deep CNN. (a) Test accuracy versus training loss. (b) The corresponding load curves during training. In the experiments, we introduce $5\%$ at epoch $20\times 200$ and $30\%$ dropout at epoch $40\times 200$.
Acknowledgments—
We are grateful to Hangfeng He, Weijie Su and Qing Qu for inspiring discussions about data separation. Shi and Dokmanić were supported by the ERC Starting Grant 852821—SWING. Pan acknowledges support from National Natural Science Foundation of China (NSFC), Grant No. 62006122 and 42230406.
@@
Appendix A S1. A folding ruler experiment
We first describe an experiment where a different cascaded mechanical system, a folding ruler, exhibits a phenomenology which is in some ways reminiscent of feature learning dynamics in DNNs. We emphasize that this reminiscence is anecdotal but nonetheless mention it since it motivated our work. The goal of the experiment is to show that noise can renegotiate the imbalance caused by friction. As shown in FIG. S1, we pin the left end of the folding ruler and pull the right end by hand. Due to friction, if we pull it very slowly and steadily, the outer layer extends far while the inner layers are close to stuck. This reminds us of lazy training where the outer layers take the largest proportion of the load. Conversely, shaking while pulling helps “activate” the inner layers and redistribute the force, and ultimately results in a uniformly distributed extension of each layer. Videos can be found at https://github.com/DaDaCheng/DNN_Spring
@@
Appendix B S2. Additional experimental results on various networks and datasets
@@
Similarly to MLPs, deep convolutional neural networks (CNNs) also learn features and progressively collapse data. Here we explore how the observation that a uniform load distribution correlates with better generalization may help guide CNN training. In FIG. 6 of the main text, we illustrate an experiment where we first train a CNN using Adam [66] at a learning rate of $10^{-4}$ on MNIST. We apply both pooling and upsampling with the same ratio after each activation function to ensure that all intermediate features have the same size. Training converges approximately after $20\times 200$ epochs, resulting in a concave load curve (purple). Next, we introduce a $5\%$ dropout which causes the training to resume and the load curve to become linear (blue). Importantly, it also improves accuracy, as predicted by the theory. Stronger noise (a $30\%$ dropout at $40\times 200$ epochs) at once results in a convex load curve (green curve) and worse generalization.
In FIG. S2, we illustrate the in-class means of the $\ell$-th mid-feature $(\bar{\bm{{x}}}_{\ell}^{k})$ at epochs 4000, 8000, and 12000. The corresponding load curves at these times are concave, linear, and convex, respectively, as shown in FIG. 6.
Figure S2: In-class means of the mid layer features in deep CNN whose training dynamics is showed in FIG. 6. We plot first channel at each layer for digits 0, 4, and 9 at epochs 4000 (concave load curve), 8000 (linear load curve), and 12000 (convex load curve).
Depth and “dead” layers—
In FIG. S3, we vary the depth of the MLP while keeping all other hyperparameters fixed. We observe that when the network is very deep, for example with $L=20$, the separation does not go further, and even degrades, as data passes through the 14th to 18th layers. This suggests that these layers are inactive (hence “dead”) and do not contribute to the overall task. Therefore, they can potentially be pruned without adversely affecting network performance. In this experiment for $L=20$, we pruned the “dead” trained layer for $\ell>13$ and added a new final linear layer, which was then retrained. This resulted in a test accuracy of 89.6%. When we retrained all layers along with the newly added final linear layer, the test accuracy increased to 90.1%.
Figure S3: Load curve with different depths $L$ in ReLU MLPs.
@@
This work primarily investigates the concavity and convexity of the training load curve. Notably, the test load curve exhibits behavior similar to that of the training load curve. In FIG. S4 we show the load curves for both the training samples (solid line) and test samples (dashed line) on the MNIST dataset. We can see that with a small learning rate, both curves are concave. Increasing the learning rate makes them both linear, and eventually convex. The most linear curve corresponds to the highest test accuracy. The same plot as FIG. 1 but with the test load curve is shown in FIG. S20, and both train and test load curves are shown together in FIG. S19.
@@
In FIG. S5, we report the coefficient of determination ($R^{2}$) to assess deviations from linearity of the load curve in FIG. 1 in the main text. We observe that the highest accuracy is always accompanied by the highest $R^{2}$ score. In all four phase diagrams shown in FIG. 1, the best test accuracy occurs around the second or third row ($\alpha=0.2$ or $\alpha=0.4$). Therefore, we compare the test accuracy with the $R^{2}$ coefficient of determination regression score in FIG. S5.
Figure S5: Comparison between test accuracy (left axis, black) and the $R^{2}$ coefficient of determination regression score (right axis, red). The data correspond to the row which contains the highest test accuracy for each phase diagram in FIG. 1 ($\alpha=0.4$ for data noise and learning rate, and $\alpha=0.2$ for dropout and batch size).
The phenomenon that the highest accuracy occurs when the load curve is linear is universally observed across a wide range of parameters and datasets. It is observed, for example, for different depths in FIG. S6 and different widths in FIG. S7.
@@
It also occurs with other datasets and optimizers. We experiment with the FashionMNIST dataset in FIG. S8 with the same default setting in (a) and a different optimizer in (b); further experiments on CIFAR10 (flatten) are shown in FIG. S9. The original color images are first resampled to a size of $3\times 10\times 10$ for computational reasons and then vectorized as input to an MLP. We note that $D_{0}$ (the data separation metric in the input layer) slightly deviates from linearity, likely due to the additional complexity of CIFAR10, but the subsequent layers closely follow the regular phenomenology discussed in the main text (FIG. S9(b)).
@@
As shown in the experiments in FIG. S9(b), noise vs. nonlinearity phase diagram remains approximately valid even when the data becomes more complex. A small deviation from this occurs when dropout is replaced by learning rate on CIFAR10: in this case, the best generalization corresponds to slightly concave load curves (FIG. S10). Since learning rate is qualitatively very different from standard “noise”, we expect that at some point it exhibits a refined phenomenology; this particular example is an intriguing avenue for further exploration.
@@
FIG. 1 and all these phase diagrams above are trained on a small portion of the dataset mainly due to computational cost. In FIG.S11, we trained on the same MLP on the full MNIST dataset; as these experiments show, the results are consistent with the results obtained on the downsampled dataset.
Figure S11: Phase diagram for dropout vs. non-linearity. The red curve shows the load curve for $\ell$ vs. $D_{\ell}$ for $\ell=1,\dots,7$. The background color represents the test accuracy. The results are obtained by training on the entire 60,000 training samples in the MNIST dataset.
We also train the MLP on the entire FashionMNIST dataset in FIG. S12. In FIG. S12(a), we fixed non-linearity and only varied batch size, while in FIG. S12, we used full batch training with learning rate as $0.0001$. In both cases, the best accuracies align with the flattest load curve.
Figure S12: Load curve of a 8-layer and 400-width MLP on trained on entire FashionMNIST dataset. (a) Variation with different batch sizes (BS) using LeakyReLU ($\alpha=0.4$) activation, where ACC denotes test accuracy. (b) Variation with LeakyReLU activation with full batch for different negative slope $\alpha$.
Additionally, we observe that as the training dataset size increases, it is preferable to use a wider MLP. With larger datasets, the load curve is more likely to become wavy or twisting (neither concave nor convex). Furthermore, we find that dropout is more effective in flattening these wavy curves compared to other noise-based methods.
Appendix C S3. Details of numerical experiments and reproducibility
@@
In Figs. 1, 4, S3, S4, S6, S7, S8, S9, S10, S11, S12, S13, S14, S15 and S16, the networks are $8$-layer fully connected MLPs (7 hidden layers), with layer width equal to $100$ unless stated otherwise. We use ReLU activations, BatchNorm in each layer, and no dropout as the default setting. All parameters are initialized using the default settings in PyTorch. Unless otherwise specified, the networks are trained using the ADAM optimizer on 2560 training samples from the MNIST dataset, with cross-entropy loss for classification tasks. The learning rate is set to be 0.001 and the batch size is 2560 unless otherwise stated. In the phase diagrams, the load curves and test accuracies are measured on the full 10,000-sample test dataset after 100 epochs of training for the experiments on data noise, dropout, and batch size, and after 200 epochs for the learning rate experiment. The results in all phase diagrams are averaged over 10 independent runs. Notably, when we replace BatchNorm with the scaling constant $c_{\ell}$ mentioned in the main text, we still observe similar convex–concave patterns in the phase diagram corresponding to noise and nonlinearity. However, without the adaptive BatchNorm, the load curves exhibit more fluctuations and are less smooth, and the variations between independent runs become more pronounced.
@@
In the left column of FIG. 4, we train the described DNN with a learning rate of 0.0001 and the batch size of 200 in panel (b); learning rate of 0.001, dropout of 0.1 and batch size of 100 in panel (c); learning rate of 0.003, and batch size of 50, and dropout of 0.2 in panel (d). In the right column, we set $k=1,y=1$ for all three cases, and $\mu_{\rightarrow}=0.11,\mu_{\leftarrow}=0.03$ and $\varepsilon=0.006$ for (b); $\mu_{\rightarrow}=\mu_{\leftarrow}=0.04$ and $\varepsilon=0.01$ for (c); $\mu_{\rightarrow}=0.04,\mu_{\leftarrow}=0.12$ and $\varepsilon=0.011$ for (d). These parameters are chosen to produce qualitatively similar curves with the previous three characteristic training dynamics. For the spring experiments in FIG. 4, we use the same noise for all blocks $\xi_{\ell}(t)=\xi_{\ell^{\prime}}(t)$, to mimic the training of DNNs in which the randomness (e.g., data noise, learning rate, and batch size) in each layer is not independent. This synchronous noise results in more similar dynamics over epochs (in particular, the fluctuations) but ultimately leads to similar load curves as independent noise, as shown in FIG. 2.
@@
In FIG. 5(a), we adopt the setting from [41]. We use SGD (instead of ADAM) with a learning rate of 0.001 to train a linear DNN. All weight matrices are initialized as random orthogonal matrices; the data is also a random orthogonal matrix with random binary labels. There is no batch normalization. We set the learning rate to $0.01$ to generate the “large step size” result (blue curve) and apply $10\%$ dropout to generate the “dropout” results (green curve).
@@
In FIG. 6, we consider a CNN with 16 channels and 6 convolutional layers on the same MNIST dataset. We use ReLU activations and BatchNorm between the convolutional layers. Pooling and upsampling are applied after each activation in such a way that each intermediate layer has the same shape, and a linear layer is applied after the final convolutional layer. The learning rate is set as 0.0001. In FIG. 5(c), the CNN has 7 convolutional layers and 20 channels for each layer, and we train the network for 2000 epoches on the CIFAR10 dataset. The other hyperparameters are the same as in the experiments in FIG. 6.
@@
Appendix D S4. Some other metrics for data separation of intermediate features
@@
$\hat{D}_{\ell}:=\log\left(\mathrm{Tr}\left(\Sigma_{\ell}^{\mathrm{w}}{\Sigma_{\ell}^{\mathrm{b}}}^{+}\right)\right),$ (S1)
@@
Figure S13: Phase diagram of the DNN load curve on the MNIST for nonlinearity vs. dropout with the data separation as $\hat{D}_{\ell}$ in Eq. (S1).
The behavior of the two metrics shows clear differences in the case of linear networks where $\hat{D}_{\ell}$ cannot capture the differences in features across different layers (FIG. S14 (c) and (d)).
Figure S14: Comparison of data separation metrics $D_{\ell}$ (Eq. (3)) in (a, c) and $\hat{D}_{\ell}$ (Eq. (S1)) in (b, d). The top two plots (a, b) show results for the ReLU DNN, while the bottom two plots (c, d) show results for the linear DNN. The curves in all four plots represent the same DNNs trained on MNIST with different dropout (dp) rates, as indicated in the legend in (c).
Appendix E S5. Regression
@@
$D_{\ell}:=\sqrt{\min_{\bm{{w}},\bm{{b}}}\frac{1}{2N}\sum_{i=1}^{N}\|\bm{{y}}_{i}-\bm{{w}}^{\top}(\bm{{x}}_{i})_{\ell}+\bm{{b}}\|^{2}_{2}},$
where $N$ is the number of data pairs and the summation is taken over all data. As shown in FIG. S15, noisier training results in a convex load curve, whereas lazy training results in a concave curve. Furthermore, the straighter line shows better generalization.
Figure S15: Regression load curves on the diabetes dataset [67] (a) and superconductivity critical temperature [68] (b) for a 7-hidden-layer ReLU network. In (a), dp denotes the dropout ratio; in (b), lr represents the learning rate; and e indicates the test RMSE.
We also test the MNIST dataset in a regression setting by using one-hot labels. The regression load curves consistently align with the phenomenon observed in the classification task in FIG. S16.
Figure S16: Regression load curves on MNIST dataset trained with one-hot label and MSE loss for a 7-hidden-layer ReLU/LeakyReLU network. In (a), dp denotes the dropout ratio and MLPs have ReLU activation function; in (b), $\alpha$ represents the native slope of LeakyReLU; and e indicates the test RMSE.
Appendix F S6. Additional details about the spring block systems
F.1 Friction in the second order system
@@
F.2 Noisy equilibiria and separation of time scales
@@
F.3 Second order Langevin equation
We can obtain a more standard Langevin dynamics formulation of Eq. (S2) by adding noise to the velocity independently of the friction, i.e., computing the friction before adding noise. We note that this is less realistic from a physical point of view. The position of the $\ell$th block $x_{\ell}=\sum_{i=1}^{\ell}d_{i}$ then obeys the equation of motion $\mathrm{d}x=v\mathrm{d}t$ with
$\mathrm{d}v_{\ell}=\left(k(\bm{L}x)_{\ell}-f(v_{\ell})-\gamma v_{\ell}\right)\mathrm{d}t+\varepsilon(t)\mathrm{d}W_{\ell}(t),$ (S4)